OpenStudy (anonymous):

It is given that f(x) = 1/x^3 − x^3, for x > 0. Show that f is a decreasing function.

OpenStudy (raden):

hint : f ' (x) < 0 can u determine f ' (x), first ?

OpenStudy (raden):

with f ' (x) is the first derivative of f(x)

OpenStudy (raden):

@accio.firebolt

OpenStudy (anonymous):

(-3x^-4) - 3x^2

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

yup, that's right or it can modif be : -3/x^4 - 3x^2, right ?

OpenStudy (raden):

or -3(1/x^4 + x^2), right ?

OpenStudy (anonymous):

@RadEn so because it's negative, then it's less than 0 so it's a decreasing function?

OpenStudy (raden):

yes, because the terms in bracket absolutly always be a real positive numbers) -3 multiplied by a positive number = negative number, in other words always < 0 " that's why f is a decreasing function "

OpenStudy (anonymous):

@RadEn thanks so much! is okay if i ask you another question?

OpenStudy (raden):

if i can i wll :)

OpenStudy (anonymous):

dy/dx = 2(3x+4)^1.5 -6x-8 It is now given that the stationary point on the curve has coordinates (-1.5). Find the equation of the curve. So i tried to find y and i got 4/15 ( 3x +4)^2.5 -3x^2 - 8x + c Then when i tried to put the coordinates in. the c i got was different from the real answer.

OpenStudy (anonymous):

please answer my question

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

sorry the conection is lagging :( hmm.. ur integration is okay yeah, for x=-1.5 gives f(x) becomes not real

OpenStudy (raden):

are u sure the point given is x = -1.5 ? it does not exist for the curve of f(x)

OpenStudy (anonymous):

the x is -1 and y is 5 for the stationary point

OpenStudy (anonymous):

\[{dy \over dx} = 2(3x + 4)^{3 \over 2} -6x -8\] \[\int \frac{dy}{dx}dx = 2\int \left((3x + 4)^{3 \over 2}-3x -4\right)dx\] \[y=\frac{4}{15}\left(3x + 4\right)^2\sqrt{3x + 4} - 3x^2 - 8x + c\] Given that x = -1 and y = 5, \[\Large 5 = \frac{4}{15}\left(1\right)^{\frac{5}{2}}-3(-1)^2+8+c\] \[5 = \frac{4}{15} + 5 + c\] \[c = -\frac{4}{15}\]

OpenStudy (anonymous):

@meepi i also got -4/15 for c but the answer is apparently 5

OpenStudy (anonymous):

if it's 5, then substituting in -1 gives 154/15 :/ It can't be 5, I'd assume the book is wrong or that you could have interpreted the problem incorrectly

OpenStudy (anonymous):

okay thanks anyway :)